OpenStudy (anonymous):

A total of 396 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold . How many adult tickets were sold?

OpenStudy (anonymous):

Systems of equations: x = number of adult tickets sold y = number of student tickets sold x + y = 396 y = 2x if y = 2x that means we can literally put 2x in for y in the first equation. x + (2x) = 396 3x = 396 x = 132 x + y = 396 132 + y = 396 y = 264

OpenStudy (anonymous):

You just did a systems but left out a step... I used substitution, which is just what you did.